# Beyond Pairwise: Global Zero-shot Temporal Graph Generation

### v2

@@ Appendix A Experimental Details
Appendix B Additional Tested LLMs
Beyond our main experiments with GPT-4o and DeepSeek-R1, we also evaluated our model using GPT-o3-mini, which achieved promising results, with an F1 score of 55.5 on NT-6 and 78.1 on OmniTemp. Additionally, we tested our method with several contemporary models accessed via the Together.ai and Google Gemini APIs. These include LLaMA (v3.1 405B, v3.3 70B) and Gemini (Flash-2.0, Pro-1.5). Our findings suggest that while all of these models [^7] Except Gemini Pro-1.5, which truncated the generation with the message: “rest of the pairs are similar, and the logic should follow the timeline explanation.” are capable of generating a complete set of temporal relations in a single step, their performance is significantly lower compared to GPT-4o, DeepSeek-R1, and GPT-o3-mini. This indicates that our method for generating complete temporal graphs currently yields the best results with more advanced models.
@@ C.3 Further Details on Event Count
Model MATRES TB-Dense NT-6 OmniTemp
@@ Appendix E OmniTemp Annotation Proccess
For the annotation process of OmniTemp (detailed in §3.1), we hired three annotators (two males and one female), all non-expert native English speakers and either undergraduate or graduate students, their location will be revealed once anonymity requirements are lifted. We instruct annotators to follow the MATRES annotation guidelines, considering only “actual” events (e.g., they won the game). Events that are “non-actual”, such as intentional, negated, recurring, conditional, or wishful (e.g., I wish they win the game), are excluded from annotation. Additionally, only the starting time of events is considered when establishing temporal relations.
The actual annotation was done on 30 news summaries, each containing approximately 500 words. The annotators used the EventFull annotation tool Eirew et al. (2025), with all events in each document already highlighted. These events were extracted using the event detection method proposed by Cattan et al. (2021), which identifies all types of events (actual and non-actual) and extracts an average of 60 event mentions per document, forming the initial set of events. We follow the same annotation protocol as proposed in EventFull. [^8] The complete annotation guidelines are available within the EventFull annotation tool and GitHub repository https://github.com/AlonEirew/EventFull. First, the annotation process begins with the selection of 15 to 18 of the most salient “actual” events from each story, following Eirew et al. (2025) which found that beyond 18 events, annotation becomes challenging for non-expert annotators. This event reduction aligns with previous efforts to decrease annotation workload by limiting the number of events considered Chambers et al. (2014); Ning et al. (2018b); Tan et al. (2024). After selecting these events, each document was annotated for temporal relations (before, after, equal, or vague) by all three annotators. Finally, majority voting was used to determine the final relation, and in cases of disagreement, the relation was labeled as vague. Further details about the annotators, time and cost are provided in Appendix E.
The total annotation time for OmniTemp, including onboarding, amounted to 85 hours, with each worker paid \$$\$15$ per hour (which is considered a fair market value in their region).
Appendix F Filling Transitive Relations
As discussed in §3.2, to assess the coverage achievable by inferring transitive relations in resources annotated only with consecutive sentences, we extracted from NarrativeTime only the relations between event pairs in consecutive sentences. We then applied a transitive closure algorithm Warshall (1962) to construct additional relations and compared the results with the original set of relations. Figure 11 presents the experimental results.
Appendix G Formal Description of GlobalConsistency
@@
Appendix H Dataset Licenses and Sources
@@
Appendix I Adjustments to the NarrativeTime Dataset
@@
Appendix J Additional Experiment Tables and Figures
Figure 7 presents the relation-wise performance of GPT-4o, analogous to the results shown for DeepSeek-R1 in Figure 4. Figure 6 presents the model performance across different relation subsets, analogous to the results shown for DeepSeek-R1 in Figure 3. Table 7 presents a comparison between common datasets used for evaluating models on the temporal relation task alongside OmniTemp. Table 6 presents the split statistics of these datasets. Figure 9 presents an example of the ZSL-Global prompt. Figure 10 presents an example of the generated timeline using the ZSL-Timeline approach.
@@
MATRES TB-Dense TCR TDD-Manual NarrativeTime OmniTemp

### v3

@@ 8 Conclusion
Acknowledgments
This work was supported by the Israel Science Foundation (grant no. 2827/21), and by funding from the Israeli Planning and Budgeting Committee (PBC).
We also used AI-based assistance (ChatGPT) for grammar and spelling corrections during manuscript preparation.
@@ Appendix A Experimental Details
Model
NT-6
OmniTemp
DeepSeek-R1 64.1 79.2
DeepSeek-V3 55.1 74.4
GPT-o3-mini 55.5 78.1
GPT-4o 58.4 73.6
Llama-3.1 405B Instruct 42.6 57.9
Llama 3.3 70B-Instruct 40.7 44.3
Gemini-flash 2.0 32.1 47.3
Table 5: Additional results of our GlobalConsistency approach when applying different LLMs, evaluated on the OmniTemp and NT-6 datasets.
GPT-4o
DeepSeek-R1
GlobalConsistency 71.2 64.9
Table 6: Results on the validation set of MAVEN-ERE Wang et al. (2022) on the sub-set of relations we selected.
Appendix B Additional Experiments
B.1 Additional Tested LLMs
Beyond our main experiments with GPT-4o and DeepSeek-R1, we also evaluated our model with additional LLMs, summarized in Table 5 together with GPT-4o and DeepSeek-R1 for ease of comparison. GPT-o3-mini and DeepSeek-V3 achieved promising results on both NT-6 (55.1 and 55.5) and OmniTemp (74.4 and 78.1). Additionally, we tested our method with several contemporary models accessed via the Together.ai and Google Gemini APIs, including LLaMA (v3.1 405B, v3.3 70B) and Gemini (Flash-2.0, Pro-1.5). Our findings suggest that while all of these models (except Gemini Pro-1.5, which truncated the generation with the message: “rest of the pairs are similar, and the logic should follow the timeline explanation.”) are capable of generating a complete set of temporal relations in a single step, they achieved much lower results. This indicates that our method for generating complete temporal graphs currently performs best with more advanced models.
B.2 Additional Tested Dataset
We conducted an additional experiment with our GlobalConsistency method on the MAVEN-ERE Wang et al. (2022) dataset, which introduces an additional domain, as it was curated from Wikipedia (in contrast to the newswire datasets used in our main experiments). MAVEN-ERE includes multiple relation categories, such as temporal, coreference, causal, and sub-events. In our setting, we used only the temporal relations portion of the dataset. For temporal relations, MAVEN-ERE defines six relation types: BEFORE, CONTAINS, OVERLAP, BEGINS-ON, ENDS-ON, and SIMULTANEOUS. To manage costs, we ran an experiment on the validation set, selecting documents with fewer than 200 pairs and considering only the BEFORE and SIMULTANEOUS relations. We applied the four-relations instruction (before, after, equal, and vague in our setting). The results are shown in Table 6, and we observe that the GlobalConsistency method achieves results comparable to those reported on the four-relation news datasets, further confirming the method’s transferability across domains.
@@ C.3 Further Details on Event Count
Model
MATRES
TB-Dense
NT-6
OmniTemp
@@ Appendix E OmniTemp Annotation Proccess
For the annotation process of OmniTemp (detailed in §3.1), we hired three annotators (two males and one female), all non-expert native English speakers and either undergraduate or graduate students. We instruct annotators to follow the MATRES annotation guidelines, considering only “actual” events (e.g., they won the game). Events that are “non-actual”, such as intentional, negated, recurring, conditional, or wishful (e.g., I wish they win the game), are excluded from annotation. Additionally, only the starting time of events is considered when establishing temporal relations.
The actual annotation was done on 30 news summaries, each containing approximately 500 words. The annotators used the EventFull annotation tool Eirew et al. (2025), with all events in each document already highlighted. These events were extracted using the event detection method proposed by Cattan et al. (2021), which identifies all types of events (actual and non-actual) and extracts an average of 60 event mentions per document, forming the initial set of events. We follow the same annotation protocol as proposed in EventFull. First, the annotation process begins with the selection of 15 to 18 of the most salient “actual” events from each story, following Eirew et al. (2025) which found that beyond 18 events, annotation becomes challenging for non-expert annotators. This event reduction aligns with previous efforts to decrease annotation workload by limiting the number of events considered Chambers et al. (2014); Ning et al. (2018b); Tan et al. (2024). After selecting these events, each document was annotated for temporal relations (before, after, equal, or vague) by all three annotators. Finally, majority voting was used to determine the final relation, and in cases of disagreement, the relation was labeled as vague.
Finally, the total annotation time for OmniTemp, including onboarding, amounted to 85 hours, with each worker paid $\mathdollar 15$ per hour (which is considered a fair market value in their region).
Appendix F Formal Description of GlobalConsistency
@@
Appendix G Dataset Licenses and Sources
@@
Appendix H Adjustments to the NarrativeTime Dataset
@@
Appendix I Additional Experiment Tables and Figures
Figure 7 presents the relation-wise performance of GPT-4o, analogous to the results shown for DeepSeek-R1 in Figure 4. Figure 6 presents the model performance across different relation subsets, analogous to the results shown for DeepSeek-R1 in Figure 3. Table 9 presents a comparison between common datasets used for evaluating models on the temporal relation task alongside OmniTemp. Table 8 presents the split statistics of these datasets. Figure 9 presents an example of the ZSL-Global prompt. Figure 10 presents an example of the generated timeline using the ZSL-Timeline approach. Figure 11 presents the experimental results for filling transitive relations in a dataset containing only temporal relations between events up to one sentence apart (similar to MATRES and TB-Dense).
@@ Appendix I Additional Experiment Tables and Figures
MATRES
TB-Dense
TCR
TDD-Manual
NarrativeTime
OmniTemp